OpenStudy (anonymous):

what is 66 2/3% of 39?

OpenStudy (sepeario):

What is that percentage as a fraction?

OpenStudy (anonymous):

again, turn it into a fraction. (66 2/3)/100 *39. = 3900/(66 + 2/3) =58.5

OpenStudy (sepeario):

Yes. Essentially what he did was multiply the fraction by the number; \[\frac{ 66 \frac{ 2 }{ 3 } }{ 100 } *\frac{ 39 }{ 1 }\]

OpenStudy (bee_see):

Hmm, I got something different. You can do it two different ways...